# Stochastic Data-Driven Predictive Control: Chance-Constraint Satisfaction with Identified Multi-step Predictors

### v1

@@ Abstract
We propose a novel data-driven stochastic model predictive control framework for uncertain linear systems with noisy output measurements. Our approach leverages multi-step predictors to efficiently propagate uncertainty, ensuring chance constraint satisfaction with minimal conservatism. In particular, we present a strategy to identify multi-step predictors and quantify the associated uncertainty using a surrogate (data-driven) state space model. Then, we utilize the derived distribution to formulate a constraint tightening that ensures chance constraint satisfaction. A numerical example highlights the reduced conservatism of handling parametric uncertainty in the proposed method compared to state-of-the-art solutions.
@@ I-A Related Work
State-space models: One of the most natural approaches to address this problem is the identification of a state-space model and the subsequent design of an MPC formulation that robustly accounts for parametric uncertainty, see, e.g.,  [9, 10, 11]. These methods typically ensure satisfaction of safety-critical constraints by using an efficient sequential propagation of the uncertainty. However, this sequential propagation often leads to overly conservative solutions.
@@
Alternatively, indirect data-driven methods first identify multi-step predictors from data [8, 7, 17, 18]. Notably, the strategy of parameterizing multi-step transitions as distinct models has already been explored in seminal works [19, 20] in the form of finite impulse response models. Nevertheless, the identification of general multi-step predictor models and their subsequent use for robust/stochastic predictive control purposes pose substantial challenges. In case of bounded noise and disturbances, set-membership identification can be leveraged to characterize multi-step predictors from data [7]. Given such bounds on multi-step predictors, robust control designs can be directly leveraged to ensure satisfaction of constraint for worst-case uncertainty [17, 7, 18]. However, such robust methods can lead to overly conservative results and in general are not suitable for rare outlier noise. In contrast, [8] proposes a strategy to identify multi-step predictors in case of unbounded stochastic disturbances with noisy state measurements. Additionally, a constraint tightening approach is introduced to ensure satisfaction of chance constraints, employing ellipsoidal confidence bounds for the model parameters. Finally, [21] presents a strategy for identifying multi-step predictors for stochastic systems and introduces a constraint tightening method. However, the approach neglects parametric uncertainty in the propagation of state covariance. Overall, ensuring chance-constraints in data-driven predictive control with stochastic output measurements remains an open problem.
@@ I-B Contribution and Outline
The primary contribution of this paper is the development of a data-driven predictive control framework that employs multi-step predictors to ensure chance-constraint satisfaction for linear stochastic systems using only input-output data. First, we define the considered stochastic optimal control problem in Section II. In Section III, we outline the proposed data-driven solution. Specifically, we provide a novel method for identifying multi-step predictors and quantifying the associated uncertainty, which will be discussed in Section IV. This method applies maximum likelihood estimation based on the innovation form of the system obtained through Kalman filter recursions of an estimated state-space model. Afterwards in Section V, the proposed predictive control problem ensures chance constraint satisfaction by accounting for the uncertainty of the estimated parameters. Overall, the proposed solution offers an alternative strategy to propagate the epistemic uncertainty of data-driven state-space models. Finally, in Section VI, we validate our approach with a numerical example, demonstrating reduced conservatism compared to state-of-the-art solutions, and conclude the paper in Section VII.
@@ I-C Notation
The set of integers within the interval $[a,\ b]\subseteq{\mathbb{R}}$ is represented by $\mathbb{I}_{[a,b]}$. For a sequence of matrices $X_{t}\in{\mathbb{R}}^{n\times m}$, with $t\in\mathbb{I}_{[t_{1},t_{2}]}$ and $t_{1},t_{2}\in\mathbb{I}_{[\geq 0]}$, the stacked matrix is denoted by $X_{[t_{1}:t_{2}]}\in{\mathbb{R}}^{n(t_{2}+1-t_{1})\times m}$. We use $I$ to denote identity matrix of appropriate dimensions. For a square matrix $A\in{\mathbb{R}}^{n\times n}$, we use $A^{1/2}$ to denote the symmetric square-root. The expected value of a random variable $x$ and the probability of an event $A$ are denoted by ${\mathbb{E}}[x]$ and $\Pr[A]$, respectively. For conditional probability of $A$ given event $B$ we use $\Pr[A\mid B]$. A variable $x$ following a Gaussian distribution with mean $\mu$ and variance $\Sigma$ is denoted as $x\sim\mathcal{N}(\mu,\Sigma)$. The quantile function of the chi-squared distribution with $n$ degrees of freedom at probability $p\in[0,1]$ is denoted by $\chi^{2}_{n}(p)$. The positive semi-definiteness of a matrix $Q$ is denoted by $Q\succeq 0$, and we write $\|x\|_{Q}^{2}=x^{\top}Qx$ for $Q\succeq 0$. The Kronecker product is represented by $\otimes$ and $\mathrm{vec}(A)\in\mathbb{R}^{nm}$ denote the operation that transforms a matrix $A\in\mathbb{R}^{n\times m}$ into a vector by sequentially stacking its columns. To denote a block-diagonal matrix with matrices $A$ and $B$ on the diagonal, we use $\mathrm{diag}(A,B)$. Similarly, to express a block-diagonal matrix with matrix $A$ repeated $n$ times on its diagonal, we use $\mathrm{diag}_{n}(A)$.
@@ II Problem Setup
with state $x_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$, control input $u_{t}\in\mathbb{R}^{n_{\mathrm{u}}}$, measured output $y_{t}\in\mathbb{R}^{n_{\mathrm{y}}}$, time $t\in\mathbb{I}_{[\geq 0]}$, disturbance $w_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$, and measurement noise $v_{t}\in\mathbb{R}^{n_{\mathrm{y}}}$. Both $w_{t}$, $v_{t}$ are independently and identically (i.i.d.) Gaussian distributed with zero mean and covariance matrices $I$, $R\succeq 0$, correspondingly. We consider the following stochastic optimal control problem:
@@ III-B Proposed Solution
Algorithm 1 Proposed Solution
1:Input: Input-output trajectory $U_{T}$, $Y_{T}$
2:User chosen probabilities $\delta\in(p,1)$, $\epsilon\in(\delta,1)$.
3:Offline Design:
4:Estimate state-space model, e.g., using Expectation-Maximization or Prediction Error Method [22, 23].
5:Estimate states and innovation covariances using Kalman filter with estimated state-space model.
6:Estimate multi-step predictors $\hat{G}_{0,k}$, $\hat{G}_{u,k}$ $k\in\mathbb{I}_{[1,N]}$ and quantify uncertainty using MLE (Prop. 1, Asm. 1).
7:Evaluate percentile function of generalized chi-squared distribution (17).
8:Online Phase:
9:Solve Problem (16) and apply control inputs $u_{[0:N]}$.
However, direct identification of the long-horizon relationships in multi-step predictors is more challenging than traditional state-space models due to correlations and larger parameter spaces. To address this, we propose an identification and uncertainty quantification scheme that leverages the Kalman filter applied with estimated state-space matrices as an intermediate step (Sec. IV). With these data-driven multi-step predictors and their associated uncertainties, we develop a predictive control strategy to ensures satisfaction of chance constraints in a minimally conservative manner (Sec. V), avoiding the limitations of sequential propagation in state-space models. The overall approach is summarized in Algorithm 1.
@@ IV Parameter Identification for Multi-step Predictors
with innovation $e_{t}\sim\mathcal{N}(0,S)$ i.i.d., (stationary) Kalman gain $L$, innovation covariance $S$, and posterior mean $x_{t\mid t}={\mathbb{E}}[x_{t}\mid y_{[1:t]}]$, all of which can be computed with Kalman filter recursions [24]. For simplicity of exposition, we consider steady-state Kalman gain innovation covariance. The following proposition provides the maximum likelihood estimate (MLE) for the multi-step predictors, assuming that the Kalman filter  (7) is available.
@@
Thus, the likelihood of the data is given by:
@@
$\propto$ $\exp\left[-\frac{1}{2}\|y_{[0:T-k]}-\Phi_{k,t}\theta_{k}\|_{\Sigma_{\mathrm{e},k}^{-1}}\right]$
${{=}}$ $\exp\left[-\frac{1}{2}\|\theta_{k}-\hat{\theta}_{k}\|_{\Sigma_{\theta,k}^{-1}}\right],$
@@
Proposition 1 presents a method for identifying multi-step predictors in their estimates. This approach utilizes the innovation form of the Kalman filter to construct a standard linear regression setup. By leveraging correlation information, we use generalized least-squares estimation to obtain the MLE [23]. Consequently, the proposed strategy inherits desired properties of the MLE framework. In particular, under mild assumptions it can be shown that the estimated parameter vector $\hat{\theta}_{k}$ is a consistent estimator of the true parameter vector $\theta_{k}$ and the errors $\theta_{k}\sim\mathcal{N}(\hat{\theta}_{k},\Sigma_{\theta,k})$ as $T\rightarrow\infty$ [23].
A key requirement for applying Proposition 1 is the availability of $x_{t\mid t}$, $G_{\mathrm{e},k}$, and $S$, computed using true state-space parameters. To bypass this, we propose estimating these entities via Kalman filter recursions with consistent estimates of the state-space matrices (1), as described in Alg. 1. A consistent estimate of the state-space model can be effectively obtained with existing solutions; e.g. Expectation-Maximization or Prediction Error Method[22, 9, 23]. This approach simplifies the parameter identification process while ensuring that the error distribution $\theta_{k}-\hat{\theta}_{k}$ converges asymptotically to the same distribution as when using true system parameters [25]. Thus, leveraging the asymptotic properties of the MLE, we consider the following uncertainty characterization for the controller design:
@@
The true multi-step predictor parameters satisfy $\theta_{k}\sim\mathcal{N}(\hat{\theta}_{k},\Sigma_{\theta,k})$, $\forall k\in\mathbb{I}_{[1,N]}$. Furthermore, the covariance matrices of the measurement and process noise are over-approximated, i.e., $\hat{R}\succeq R$ and $\hat{G}_{\mathrm{w,k}}\hat{G}_{\mathrm{w,k}}^{\top}\succeq G_{\mathrm{w,k}}G_{\mathrm{w,k}}^{\top}$.
Assumption 1 establishes distributions over the unknown vectors $\theta_{k}$. For the remainder of this paper, we assume that Assumption 1 holds and we denote $G_{\mathrm{w,k}}=\hat{G}_{\mathrm{w,k}}$, $R=\hat{R}$.
@@ V Stochastic Predictive Control with data-driven Multi-Step Predictors
In the following, we reformulate Problem (2) using the uncertain multi-step predictors derived in Section IV. To jointly address the stochastic uncertainty in the model parameters $\theta_{k}$ and the future disturbances $w_{[0,N]}$, we employ the concept of robustness in probability [26, Lemma V.3]. Specifically, we construct sets $\Theta_{k,j}$ such that $\Pr[\theta_{k}\in\Theta_{k,j}]\geq\delta$, where $\delta\in(p,1)$ is a user-chosen probability. Then, we enforce the chance constraints (2b) robustly for all parameters $\theta\in\Theta_{k,j}$ with an increased probability of $\tilde{p}=p/\delta$. The following proposition recaps a simple approach to address this problem by using the standard confidence ellipsoids:
@@
$h_{j}^{\top}[\hat{G}_{0,k},\ \hat{G}_{\mathrm{u},k}]\begin{bmatrix}\bar{x}_{0}\\ u_{[0,k-1]}\end{bmatrix}+d_{\delta}\left\|{\Sigma_{\mathrm{\theta},k}^{1/2}\left(\begin{bmatrix}\bar{x}_{0}\\ u_{[0,k-1]}\end{bmatrix}\otimes h_{j}\right)}\right\|$
@@
with $d_{\delta}=\sqrt{\chi^{2}_{n_{\theta,k}}(\delta)}$, $d_{k,j}=h_{j}^{\top}(G_{\mathrm{w},k}G_{\mathrm{w},k}^{\top}+R)h_{j}$,
$f_{k,j}$ $=\max_{\theta_{k}:\|\theta_{k}-\hat{\theta}_{k}\|_{\Sigma_{\theta,k}}\leq d_{\delta}}\theta_{k}^{\top}M_{k,j}\theta_{k},$ (14)
@@
Proposition 2 first establishes confidence ellipsoids $\Theta_{k,j}$ using the provided Gaussian distribution and then derives tight upper bounds for the predicted output. However, when the parameter dimension $n_{\theta}$ is large, the resulting confidence ellipsoids become overly conservative and fail to accurately capture the worst-case parameters relevant to the specific control problem. Instead, we constructs tight sets $\Theta_{k,j}$ separately that are independent of the number of parameters. In particular, we improve the constraint tightening (2) with the following optimization problem:
@@
The optimization problem (16) is a convex second-order cone program. The following theorem establishes satisfaction of the considered chance constraints (2b):
@@
Lastly, $\theta_{k}\in\Theta_{k,j}$ implies the satisfaction of constraints (2b) with probability $\tilde{p}$ since $y_{k}\sim\mathcal{N}(\bar{y}_{k},\Sigma_{\mathrm{y},k})$. Consequently,
$\Pr[h_{j}^{\top}y_{k}\leq 1\mid\theta_{k}\in\Theta_{k,j}]\Pr[\theta_{k}\in\Theta_{k,j}]\geq\tilde{p}\delta=p.$
The proposed constraint tightening (16b) ensures satisfaction of the chance constraints with minimal conservatism. The only conservatism introduced arises from using union bound for combining uncertainty due to initial state’s mean and covariance. Notably, the tightening does not depend on the number of parameters $n_{\theta}$ directly and thus induces less conservatism compared to Prop. 2 as parameter size increases. Finally, the proposed tightening strategy can be easily extended to handle two-sided constraints by replacing the constants $c_{\epsilon}$ and $c_{\tilde{p}}$ with their two-sided counterparts [2].

### v2

@@ Abstract
We propose a novel data-driven stochastic model predictive control framework for uncertain linear systems with noisy output measurements. Our approach leverages multi-step predictors to efficiently propagate uncertainty, ensuring chance constraint satisfaction. In particular, we present a strategy to identify multi-step predictors and quantify the associated uncertainty using a surrogate (data-driven) state space model. Then, we utilize the derived distribution to formulate a constraint tightening that ensures chance constraint satisfaction despite the parametric uncertainty. A numerical example highlights the reduced conservatism of handling parametric uncertainty in the proposed method compared to state-of-the-art solutions.
Index Terms:
Predictive control for linear systems, Data driven control, Identification for control
@@ I-A Related Work
State-space models: One of the most natural approaches to address this problem is the identification of a state-space model and the subsequent design of an MPC formulation that robustly accounts for parametric uncertainty, see, e.g.,  [9, 10]. These methods typically ensure satisfaction of safety-critical constraints by using an efficient sequential propagation of the uncertainty. However, this sequential propagation often leads to overly conservative solutions.
@@
Alternatively, indirect data-driven methods first identify multi-step predictors from data [8, 7, 16, 17]. Notably, the strategy of parameterizing multi-step transitions as distinct models has already been explored in seminal works [18] in the form of finite impulse response models. Nevertheless, the identification of general multi-step predictor models and their subsequent use for robust/stochastic predictive control purposes pose substantial challenges. In case of bounded noise and disturbances, set-membership identification can be leveraged to characterize multi-step predictors from data [7]. Given such bounds on multi-step predictors, robust control designs can be directly leveraged to ensure satisfaction of constraint for worst-case uncertainty [16, 7, 17]. However, such robust methods can lead to overly conservative results and in general are not suitable for rare outlier noise. In contrast, [8] proposes a strategy to identify multi-step predictors in case of unbounded stochastic disturbances with noisy state measurements. Additionally, a constraint tightening approach is introduced to ensure satisfaction of chance constraints, employing ellipsoidal confidence bounds for the model parameters. Finally, [19] presents a strategy for identifying multi-step predictors for stochastic systems and introduces a constraint tightening method. However, the approach neglects parametric uncertainty in the propagation of state covariance. Overall, ensuring chance-constraints in data-driven predictive control with stochastic output measurements remains an open problem.
@@ I-B Contribution and Outline
The primary contribution of this paper is the development of a data-driven predictive control framework that employs multi-step predictors to ensure chance-constraint satisfaction for linear stochastic systems using only input-output data. First, we define the considered stochastic optimal control problem in Section II. In Section III, we outline the proposed data-driven solution. Specifically, we provide a novel method for identifying multi-step predictors and quantifying the associated uncertainty, which will be discussed in Section IV. This method applies Maximum Likelihood Estimation (MLE) based on the innovation form of the system obtained through Kalman filter recursions of an estimated state-space model. Afterwards in Section V, the proposed predictive control problem ensures chance constraint satisfaction by accounting for the uncertainty of the estimated parameters. Overall, the proposed solution offers an alternative strategy to propagate the epistemic uncertainty of data-driven state-space models. Finally, in Section VI, we validate our approach with a numerical example, demonstrating reduced conservatism compared to state-of-the-art solutions, and conclude the paper in Section VII.
@@ I-C Notation
The set of integers within the interval $[a,\ b]\subseteq{\mathbb{R}}$ is represented by $\mathbb{I}_{[a,b]}$. For a sequence of matrices $X_{t}\in{\mathbb{R}}^{n\times m}$, with $t\in\mathbb{I}_{[t_{1},t_{2}]}$ and $t_{1},t_{2}\in\mathbb{I}_{[\geq 0]}$, the stacked matrix is denoted by $X_{[t_{1}:t_{2}]}\in{\mathbb{R}}^{n(t_{2}+1-t_{1})\times m}$. We use $I$ to denote identity matrix of appropriate dimensions. For a square matrix $A\in{\mathbb{R}}^{n\times n}$, we use $A^{1/2}$ to denote the symmetric square-root. The expected value of a random variable $x$ and the probability of an event $A$ are denoted by ${\mathbb{E}}[x]$ and $\Pr[A]$, respectively. For conditional probability of $A$ given event $B$ we use $\Pr[A\mid B]$. A variable $x$ following a Gaussian distribution with mean $\mu$ and variance $\Sigma$ is denoted as $x\sim\mathcal{N}(\mu,\Sigma)$. The quantile function of the chi-squared distribution with $n$ degrees of freedom at probability $p\in[0,1]$ is denoted by $\chi^{2}_{n}(p)$. The positive semi-definiteness of a matrix $Q$ is denoted by $Q\succeq 0$, and we write $\|x\|_{Q}^{2}=x^{\top}Qx$ for $Q\succeq 0$. The Kronecker product is represented by $\otimes$ and $\mathrm{vec}(A)\in\mathbb{R}^{nm}$ denote the operation that transforms a matrix $A\in\mathbb{R}^{n\times m}$ into a vector by sequentially stacking its columns. To denote a block-diagonal matrix with matrices $A$ and $B$ on the diagonal, we use $\mathrm{diag}(A,B)$. Similarly, to express a block-diagonal matrix with matrix $A$ repeated $n$ times on its diagonal, we use $\mathrm{diag}_{n}(A)$. $A\propto B$ indicates direct proportion between terms $A$ and $B$.
@@ II Problem Setup
with state $x_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$, control input $u_{t}\in\mathbb{R}^{n_{\mathrm{u}}}$, measured output $y_{t}\in\mathbb{R}^{n_{\mathrm{y}}}$, time $t\in\mathbb{I}_{[\geq 0]}$, disturbance $w_{t}\in\mathbb{R}^{n_{\mathrm{w}}}$, and measurement noise $v_{t}\in\mathbb{R}^{n_{\mathrm{y}}}$. Both $w_{t}$, $v_{t}$ are independently and identically (i.i.d.) Gaussian distributed with zero mean and covariance matrices $I$, $R\succ 0$, correspondingly. We consider the following stochastic optimal control problem:
@@ III-B Proposed Solution
Algorithm 1 General approach
1:Input: Input-output trajectory
2:Offline Design:
3:Estimate state-space model, e.g., using Expectation-Maximization or Prediction Error Method [20, 21].
4:Estimate states and innovation covariances using Kalman filter with estimated state-space model.
5:Estimate multi-step predictors and quantify uncertainty using MLE.
6:Online Phase:
7:Solve deterministic optimization problem and apply inputs $u_{[0,N]}$.
Notably, direct identification of the long-horizon relationships in multi-step predictors is more challenging than traditional state-space models due to correlations and larger parameter spaces.
To address this, we propose an identification and uncertainty quantification scheme that leverages the Kalman filter applied with estimated state-space matrices as an intermediate step (Sec. IV). With these data-driven multi-step predictors and their associated uncertainties, we develop a predictive control strategy to ensures satisfaction of chance constraints in a despite the parametric uncertainty(Sec. V), avoiding the limitations of sequential propagation in state-space models. The overall approach is summarized in Algorithm 1.
@@ IV Parameter Identification for Multi-step Predictors
with innovation $e_{t}\sim\mathcal{N}(0,S)$ i.i.d., (stationary) Kalman gain $L$, innovation covariance $S$, and posterior mean $x_{t\mid t}={\mathbb{E}}[x_{t}\mid y_{[1:t]}]$, all of which can be computed with Kalman filter recursions. For simplicity of exposition, we consider steady-state Kalman gain innovation covariance. The following proposition provides the maximum likelihood estimate for the multi-step predictors, assuming that the Kalman filter  (7) is available.
@@
Given the Gaussian noise $\tilde{e}[0:T-k],k$, the likelihood of the data satisfies:
@@
(${{\propto}}$ $\exp\left[-\frac{1}{2}\|y_{[0:T-k]}-\Phi_{k,t}\theta_{k}\|^{{\color[rgb]{0,0,1}2}}_{\Sigma_{\mathrm{e},k}^{-1}}\right]$
${{=}}$ $\exp\left[-\frac{1}{2}\|\theta_{k}-\hat{\theta}_{k}\|^{{\color[rgb]{0,0,1}2}}_{\Sigma_{\theta,k}^{-1}}\right],$
@@
Proposition 1 presents a method for identifying multi-step predictors in their estimates. This approach utilizes the innovation form of the Kalman filter to construct a standard linear regression setup. By leveraging correlation information, we use generalized least-squares estimation to obtain the MLE [21]. Consequently, the proposed strategy inherits desired properties of the MLE framework. In particular, under mild assumptions it can be shown that the estimated parameter vector $\hat{\theta}_{k}$ is a asymptotically consistent estimator of the true parameter vector $\theta_{k}$ and the errors ${\color[rgb]{0,0,0}\theta_{k}-\hat{\theta}_{k}}\sim\mathcal{N}(0,\Sigma_{\theta,k})$ as $T\rightarrow\infty$ [21].
A key requirement for applying Proposition 1 is the availability of $x_{t\mid t}$, $G_{\mathrm{e},k}$, and $S$, computed using true state-space parameters. To bypass this, we propose estimating these entities via Kalman filter recursions with asymptotically consistent estimates of the state-space matrices (1), as described in Alg. 1. An asymptotically consistent estimate of the state-space model can be effectively obtained with existing solutions; e.g. Expectation-Maximization or Prediction Error Method[20, 9, 21]. This approach simplifies the parameter identification process while ensuring that the error distribution $\theta_{k}-\hat{\theta}_{k}$ converges asymptotically to the same distribution as when using true system parameters [22]. The following assumption introduces an approximation, assuming that the asymptotic Gaussian distribution of the MLE holds despite the utilization of a finite data length $T<\infty$.
@@
The multi-step predictor parameter estimates satisfy $\theta_{k}-\hat{\theta}_{k}\sim\mathcal{N}(0,\Sigma_{\theta,k})$, $\forall k\in\mathbb{I}_{[1,N]}$. Furthermore, the covariance matrices of the measurement and process noise are over-approximated, i.e., $\hat{R}\succeq R$ and $\hat{G}_{\mathrm{w,k}}\hat{G}_{\mathrm{w,k}}^{\top}\succeq G_{\mathrm{w,k}}G_{\mathrm{w,k}}^{\top}$.
This approximation provides the basis for our controller design. We note that similar approximations are common for data-driven control with stochastic noise [6, 12]. For the remainder of this paper, we assume that Assumption 1 holds and we denote $G_{\mathrm{w,k}}=\hat{G}_{\mathrm{w,k}}$, $R=\hat{R}$.
@@ V Stochastic Predictive Control with data-driven Multi-Step Predictors
In the following, we reformulate Problem (2) using the uncertain multi-step predictors derived in Section IV. To jointly address the stochastic uncertainty in the model parameters $\theta_{k}$ and the future disturbances $w_{[0,N]}$, we employ the concept of robustness in probability [23, Lemma V.3]. Specifically, we construct sets $\Theta_{k,j}$ such that $\Pr[\theta_{k}\in\Theta_{k,j}]\geq\delta$, where $\delta\in(p,1)$ is a user-chosen [^1] A natural choice is $\delta=\tilde{p}=\sqrt{p}$. probability. Then, we enforce the chance constraints (2b) robustly for all parameters $\theta\in\Theta_{k,j}$ with an increased probability of $\tilde{p}=p/\delta$. The following proposition recaps a simple approach to address this problem by using the standard confidence ellipsoids $\Theta_{k,j}$:
@@
$h_{j}^{\top}[\hat{G}_{0,k},\ \hat{G}_{\mathrm{u},k}]\begin{bmatrix}\bar{x}_{0}\\ u_{[0,k-1]}\end{bmatrix}+{\color[rgb]{0,0,0}d_{\delta,k}}\left\|{\Sigma_{\mathrm{\theta},k}^{1/2}\left(\begin{bmatrix}\bar{x}_{0}\\ u_{[0,k-1]}\end{bmatrix}\otimes h_{j}\right)}\right\|$
@@
with ${\color[rgb]{0,0,0}d_{\delta,k}}=\sqrt{\chi^{2}_{n_{\theta,k}}(\delta)}$, $d_{k,j}=h_{j}^{\top}(G_{\mathrm{w},k}G_{\mathrm{w},k}^{\top}+R)h_{j}$,
$f_{k,j}$ $=\max_{\theta_{k}:\|\theta_{k}-\hat{\theta}_{k}\|_{\Sigma_{\theta,k}}\leq{\color[rgb]{0,0,0}d_{\delta,k}}}\theta_{k}^{\top}M_{k,j}\theta_{k},$ (14)
@@
Proposition 2 first establishes confidence ellipsoids $\Theta_{k,j}$ using the provided Gaussian distribution and then derives tight upper bounds for the predicted output. However, when the parameter dimension $n_{\theta}$ is large, the resulting confidence ellipsoids become overly conservative. Instead, we constructs tight sets $\Theta_{k,j}$ separately that are independent of the number of parameters. In particular, we improve the constraint tightening (2) with the following optimization problem:
@@
The optimization problem (16) is a convex second-order cone program, where the linear constraints of the nominal problem (6) are replaced by second-order cone constraints that ensure chance constraint satisfaction under parametric uncertainty. The following theorem establishes satisfaction of the considered chance constraints (2b):
@@
Finally, $\theta_{k}\in\Theta_{k,j}$ implies the satisfaction of constraints (2b) with probability $\tilde{p}$ since $y_{k}\sim\mathcal{N}(\bar{y}_{k},\Sigma_{\mathrm{y},k})$. Consequently,
$\Pr[h_{j}^{\top}y_{k}\leq 1\mid\theta_{k}\in\Theta_{k,j}]\Pr[\theta_{k}\in\Theta_{k,j}]\geq\tilde{p}\delta=p.\qed$
The tightened constraints (16b) in the proposed control problem provide a sufficient condition for chance constraint satisfaction. The term $d_{\delta,k}$ in (2) grows with the parameter dimension $n_{\theta,k}$, increasing conservatism with a larger horizon. In contrast, using a parameter set represented by a single half-space constraint, this term is replaced by $c_{\epsilon}$ in Problem (16), which is independent of parameter dimension $n_{\theta,k}$ and horizon $k$, reducing conservatism.. The only conservatism introduced arises from using union bound for combining uncertainty due to initial state’s mean and covariance. Finally, the proposed tightening strategy can be easily extended to handle two-sided constraints by replacing the constants $c_{\epsilon}$ and $c_{\tilde{p}}$ with their two-sided counterparts [2].